an iceberg is floating in ocean.what fraction of its volume is above water(density of ice=900kg/m3 and density of ocean water-1030kg/m3)

Grade:11

1 Answers

SURESH
8 years ago
dear ariana,
                 we know the formula for density=m/v in this formula density is inversely proportional to the volume so,in your question we imagine  D1=desity of ice & D2=desity of water similarly V1 & V2 are the volumes of ice and water respectively so,D1/D2=V2/V1.